### v4

# Topological Magnon-Plasmon Hybrids

Abstract
We study magnon-plasmon coupling in effectively two-dimensional stacks of van der Waals layers in the context of the band structure topology. Invoking the quasiparticle approximation, we show that the magnetic dipole coupling between the plasmons in a metallic layer and the magnons in a neighboring magnetic layer gives rise to a Berry curvature. As a result, the hybrid quasiparticles acquire an anomalous velocity, leading to intrinsic anomalous thermal Hall and spin-Nernst effects in ferromagnets and antiferromagnets. We propose magnetic layers supporting skyrmion crystals as a platform to realize chiral magnon-plasmon edge states, inviting the notion of topological magnon-plasmonics.
Introduction.
The interplay of collective excitations in condensed matter—such as phonons, magnons, and plasmons—is a topic of interest to both fundamental science and technology. The coupling of quasiparticles leads to their hybridization [1, 2], resulting in spectral anticrossings that are explored in the context of quantum hybrid systems [3, 4], band structure quantum geometry [5, 6, 7, 8], and light-matter coupling [2]. The coupling of magnons to plasmons, in particular, may serve to control hybrid states of charge density and magnetization fluctuations, thereby interfacing the charge and spin domains in the solid state. While in three dimensions there is a large energy mismatch between magnons and the gapped plasmons in bulk [9, 10], magnon-plasmon coupling in surface magnon polaritons leads to the sought-after anticrossings [11, 12, 13, 14, 15, 16]. The discovery of two-dimensional materials that can support gapless plasmons, such as graphene [17, 18, 19], and magnetic order [20, 21, 22], has opened up a promising platform for bulk magnon-plasmon hybridization [23, 24, 25, 26, 27, 28, 29]. For instance, magnons and plasmons supported in a magnetic monolayer metal have been predicted to hybridize due to spin-orbit coupling [25, 29]. Magnon-plasmon hybrids are also proposed to arise from magnetic dipole coupling in a bilayer of graphene and an insulating ferromagnet [23]. Excitingly, topological spin-plasma waves have been predicted in a heterostructure of a three-dimensional topological insulator and an insulating ferromagnetic layer, where the hybridization is enabled by the strong spin-orbit coupling in the Dirac surface state [30]. Motivated by the existence of topological polaritons (e.g., photon-exciton [5, 31] or photon-phonon hybrids [32]), here we raise the question whether magnon-plasmon coupling in less complex and effectively two-dimensional structures can give rise to band-geometric and topological effects that could lead to signatures in transport, and invite the notion of topological magnon-plasmonics.
We show that magnon-plasmon coupling in two dimensions gives rise to hybrid quasiparticles with finite Berry curvature, which we calculate explicitly for effective bilayer structures of a metal and a magnetic insulator layer coupled by magnetic dipole coupling. In uniformly ordered magnetic materials, such as ferromagnets and antiferromagnets, we identify finite Chern numbers and anomalous thermal Hall and spin-Nernst transport responses caused by the nontrivial band geometry. In addition, in the case of a magnetic skyrmion lattice, schematically shown in Fig. 1, we study the emergence of chiral magnon-plasmon edge states. These may serve to guide hybrid states of light, charge density oscillations, and magnetic oscillations in magnonic-plasmonic networks and hybridized magnonic materials [33].
Figure 1: Sketch of topological chiral edge magnon-plasmons in a bilayer of a metal (top layer) and a magnetic system (bottom layer). The magnons are coherently coupled to the plasmons by magnetic dipole coupling, causing a topologically non-trivial spectral anticrossing in the bulk, which supports the chiral hybrid excitations at the edges of the sample (indicated by the water-wave-like feature and the yellow wave packet with an arrow).
Model. Inspired by the setup in Ref. 23, we consider a heterostructure of a metallic and a magnetic layer in the $xy$ plane (decoupled by an intermediate spacer). The Hamiltonian of the full system is given by
$H^{(i)}=H_{\text{magnet}}^{(i)}+H_{\text{plasmon}}+H_{\text{coupling}}.$ (1)
The effective spin Hamiltonian $H^{(i)}_{\text{magnet}}$ describes the magnetic system in terms of localized spin operators $\boldsymbol{S}_{i}$, with site index $i$. We provide the details below, where we will consider three cases of increasing complexity: ferromagnets ($i=1$), antiferromagnets ($i=2$), and skyrmion crystals ($i=3$). The metallic layer is assumed to support plasmons with energy $\Omega_{\boldsymbol{k}}$, created (destroyed) by $a^{\dagger}_{\boldsymbol{k}}$ ($a_{\boldsymbol{k}}$), and modeled as noninteracting bosons, so that $H_{\text{plasmon}}=\hbar\sum_{\boldsymbol{k}}\Omega_{\boldsymbol{k}}a^{\dagger}_{\boldsymbol{k}}a_{\boldsymbol{k}}$. For concreteness, we adopt parameters of graphene to estimate $\Omega_{\boldsymbol{k}}$ [34], as reviewed in the Supplemental Material (SM) [35], where we also provide material details of a thin dielectric spacer. For the perturbative coupling between the layers, we assume magnetic dipole coupling, $H_{\text{coupling}}=g\mu_{\text{B}}\sum_{i}\boldsymbol{B}_{\text{p}}(\boldsymbol{r}_{i})\cdot\boldsymbol{S}_{i}$, with g-factor $g$, Bohr’s magneton $\mu_{\text{B}}$, and the dynamic magnetic field of the plasmon at coordinate $\boldsymbol{r}_{i}$,
$\boldsymbol{B}_{\text{p}}(\boldsymbol{r}_{i})$ $=\sum_{\boldsymbol{k}}\mathcal{B}_{\boldsymbol{k}}(z)\mathrm{e}^{\mathrm{i}\boldsymbol{k}\cdot\boldsymbol{r}_{i}}\,\left(\hat{\boldsymbol{k}}\times\hat{\boldsymbol{z}}\right)\,a_{\boldsymbol{k}}+\textrm{h.c.}$ (2)
Its Fourier amplitudes $\mathcal{B}_{\boldsymbol{k}}(z)$ have been calculated previously [34, 23] and are reproduced in the SM [35]. Note that the momentum reads $\boldsymbol{k}=(k_{x},k_{y},0)^{\text{T}}$, such that $\boldsymbol{B}_{\text{p}}(\boldsymbol{r}_{i})$ lies in the $xy$ plane. Hats indicate unit vectors.
Ferromagnet.
To provide the basic intuition for the topological magnon-plasmon hybridization, we first consider a simple square-lattice ferromagnet, with $N_{z}$ layers along the $z$-direction, whose spin Hamiltonian reads as
$H^{(1)}_{\text{magnet}}=-\frac{J}{2}\sum_{\langle i,j\rangle}\boldsymbol{S}_{i}\cdot\boldsymbol{S}_{j}-K\sum_{i}\left(S_{i}^{z}\right)^{2},$ (3)
[…]
The full bilinear part of $H^{(1)}$ in Eq. (1) reads $H^{(1)}=\frac{1}{2}\sum_{\boldsymbol{k}}\boldsymbol{\phi}^{\dagger}_{\boldsymbol{k}}H_{\boldsymbol{k}}\boldsymbol{\phi}_{\boldsymbol{k}}$ where $\boldsymbol{\phi}^{\dagger}_{\boldsymbol{k}}=(b^{\dagger}_{\boldsymbol{k}},a^{\dagger}_{\boldsymbol{k}},b_{-\boldsymbol{k}},a_{-\boldsymbol{k}})$. We drop the anomalous couplings ($a^{\dagger}b^{\dagger}$ and $ab$) between magnons and plasmons, because their coupling strength is much smaller than the energy at which the bands cross, and thus they give negligible corrections to the hybridization. Thus, the dynamics of the coupled system is approximated by $H^{(1)}\approx\sum_{\boldsymbol{k}}\boldsymbol{\psi}^{\dagger}_{\boldsymbol{k}}\tilde{H}_{\boldsymbol{k}}\boldsymbol{\psi}_{\boldsymbol{k}}$, with $\boldsymbol{\psi}^{\dagger}_{\boldsymbol{k}}=(b^{\dagger}_{\boldsymbol{k}},a^{\dagger}_{\boldsymbol{k}})$ and
$\tilde{H}_{\boldsymbol{k}}=\begin{pmatrix}\hbar\omega_{\boldsymbol{k}}&G_{k}\mathrm{e}^{\mathrm{i}\varphi_{\boldsymbol{k}}}\\ G_{k}\mathrm{e}^{-\mathrm{i}\varphi_{\boldsymbol{k}}}&\hbar\Omega_{\boldsymbol{k}}\end{pmatrix},\quad\tan\varphi_{\boldsymbol{k}}=\frac{\hat{k}_{y}}{\hat{k}_{x}}.$ (4)
The coupling $G_{k}\mathrm{e}^{\mathrm{i}\varphi_{\boldsymbol{k}}}$ derives from $\boldsymbol{B}_{\text{p}}(\boldsymbol{r}_{i})\cdot\boldsymbol{S}_{i}=B^{-}_{\text{p}}(\boldsymbol{r}_{i})S^{+}_{i}+B^{+}_{\text{p}}(\boldsymbol{r}_{i})S^{-}_{i},$ with $B^{\pm}_{\text{p}}(\boldsymbol{r}_{i})=[B^{x}_{\text{p}}(\boldsymbol{r}_{i})\pm\mathrm{i}B^{y}_{\text{p}}(\boldsymbol{r}_{i})]/2\propto\mathrm{e}^{\mathrm{i}\varphi_{\boldsymbol{k}}}$. We provide the full expression of $G_{k}$ in the SM [35], but we note here that the $G_{k}\propto\sqrt{N_{z}}$, since a single plasmon mode couples to the Kittel-like mode along the $z$-direction, formed by $N_{z}$ layers.
Diagonalizing $\tilde{H}_{\boldsymbol{k}}$ yields two energies $E_{\boldsymbol{k}}^{\pm}$. As the hybridization happens at long wavelengths, we approximate $\omega_{\boldsymbol{k}}\approx\omega_{k}$ and $\Omega_{\boldsymbol{k}}\approx\Omega_{k}$, where $k=|\boldsymbol{k}|$. The remaining dependence of $H_{\boldsymbol{k}}$ on the direction of $\boldsymbol{k}$ comes from the phase factors. Thus, the eigenvectors $\boldsymbol{w}^{\pm}_{\boldsymbol{k}}$ of $\tilde{H}_{\boldsymbol{k}}$ can be written as $\boldsymbol{w}_{\boldsymbol{k}}^{\pm}=U_{\boldsymbol{k}}\boldsymbol{v}_{k}^{\pm}$, where $U_{\boldsymbol{k}}=\text{Diag}(\mathrm{e}^{\mathrm{i}\varphi_{\boldsymbol{k}}/2},\mathrm{e}^{-\mathrm{i}\varphi_{\boldsymbol{k}}/2})$ and $\boldsymbol{v}_{k}^{\pm}$ are the eigenvectors of $H_{\boldsymbol{k}=(k_{x},0)}$ for which $\varphi_{\boldsymbol{k}}=0$. The Berry curvature reduces to
$F^{\pm}_{\boldsymbol{k}}=\mathrm{i}\left(\frac{\partial\boldsymbol{w}_{\boldsymbol{k}}^{\pm,\dagger}}{\partial k_{x}}\frac{\partial\boldsymbol{w}_{\boldsymbol{k}}^{\pm}}{\partial k_{y}}-\frac{\partial\boldsymbol{w}_{\boldsymbol{k}}^{\pm,\dagger}}{\partial k_{y}}\frac{\partial\boldsymbol{w}_{\boldsymbol{k}}^{\pm}}{\partial k_{x}}\right)=-\frac{1}{2k}\frac{\partial}{\partial k}\left(\boldsymbol{v}_{k}^{\pm,\dagger}\sigma_{3}\boldsymbol{v}_{k}^{\pm}\right),$ (5)
with $\sigma_{3}=\text{Diag}(1,-1)$; it is rotationally symmetric, $F^{\pm}_{\boldsymbol{k}}=F^{\pm}_{k}$. The Chern number is obtained as
$C^{\pm}=\int\frac{F^{\pm}_{\boldsymbol{k}}}{2\pi}\,\mathrm{d}^{2}k=\frac{\left(\boldsymbol{v}_{k}^{\pm,\dagger}\sigma_{3}\boldsymbol{v}_{k}^{\pm}\right)_{k\to\infty}-\left(\boldsymbol{v}_{k}^{\pm,\dagger}\sigma_{3}\boldsymbol{v}_{k}^{\pm}\right)_{k\to 0}}{2}$ (6)
and we find a topological phase with $C^{\pm}=\pm 1$ [35]. The above considerations of topology agree with those in Ref. 37 on classically coupled spin waves and electromagnetic fields.
Figure 2: Magnon-plasmon coupling in a ferromagnet-metal bilayer. (a) Quasiparticle dispersion with color indicating magnon and plasmon character. The Chern number $C_{n}$ of the upper and lower band is indicated. (b) Thermal Hall conductivity $\kappa_{xy}$ as a function of temperature $T$. Here $\Delta=0.2\,\mathrm{meV}$ is the ferromagnetic magnon gap.
As an example, we choose $S=5/2$, $K=0.04\,\mathrm{meV}$, $J=1.15\,\mathrm{meV}$ and $N_{z}=10^{3}$ and plot the magnon-plasmon anticrossing in Fig. 2(a), with color indicating the magnon and plasmon character. The emerging Berry curvature gives rise to an anomalous velocity and causes an intrinsic thermal Hall effect, that is, a heat current density $q_{x}=-\kappa_{xy}\partial_{y}T$ directed along the $x$ direction as response to a temperature gradient $\partial_{y}T$ in the $y$ direction, with the thermal Hall conductivity $\kappa_{xy}$ given by [38, 39, 40, 41]
$\kappa_{xy}$ $=-\frac{k_{\text{B}}^{2}T}{\hbar A}\sum_{n=\pm}\sum_{\boldsymbol{k}}\left[c_{2}(\rho_{n,\boldsymbol{k}})-\frac{\pi^{2}}{3}\right]F_{\boldsymbol{k}}^{n},$ (7)
where $k_{B}$ is the Boltzmann constant, $T$ is temperature, and $A$ is the total area of the system. Furthermore, $c_{2}(x)=(1+x)(\ln{\frac{1+x}{x}})^{2}-(\ln{x})^{2}-2\text{Li}_{2}(-x)$, where $\text{Li}_{2}(x)$ is the dilogarithm function, and $\rho_{n,\boldsymbol{k}}=[\mathrm{e}^{E_{n,\boldsymbol{k}}/(k_{B}T)}-1]^{-1}$ is the Bose-Einstein function. As shown in Fig. 2(b), the thermal Hall conductivity gets thermally activated and reaches a plateau around $k_{\text{B}}T\sim\Delta$. With typical FM resonance frequencies in the range of some tens of GHz, temperatures of a few Kelvin are enough to saturate $\kappa_{xy}$. Here, we have dropped the anomalous couplings in the calculation of the thermal Hall conductivity, but we show in the SM [35] that including the anomalous couplings does not affect qualitatively the thermal Hall conductivity.
Antiferromagnet.
As a more complex model, we consider an effectively two-dimensional easy-axis antiferromagnet on the square lattice described by the spin Hamiltonian
$H_{\text{magnet}}^{(2)}=\frac{J}{2}\sum_{\langle i,j\rangle}\boldsymbol{S}_{i}\cdot\boldsymbol{S}_{j}-K\sum_{i}\left(S_{i}^{z}\right)^{2},$ (8)
where $J>0$ is antiferromagnetic nearest-neighbor exchange interaction and $K>0$ out-of-plane easy-axis anisotropy.
We perform an antiferromagnetic Holstein-Primakoff transformation [35], after which the bilinear Hamiltonian reads $H^{(2)}\approx\frac{1}{2}\sum_{\boldsymbol{k}}\boldsymbol{\Phi}^{\dagger}_{\boldsymbol{k}}H_{\boldsymbol{k}}\boldsymbol{\Phi}_{\boldsymbol{k}}$ in the basis $\boldsymbol{\Phi}^{\dagger}_{\boldsymbol{k}}=(b_{\boldsymbol{k},1}^{\dagger},b_{\boldsymbol{k},2}^{\dagger},b_{-\boldsymbol{k},1},b_{-\boldsymbol{k},2},a^{\dagger}_{\boldsymbol{k}},a_{-\boldsymbol{k}})$, where $b_{\boldsymbol{k},n}^{\dagger}$ creates a spin flip on the $n$th sublattice of the antiferromagnet ($n\in\{1,2\}$). The Hamilton matrix $H_{\boldsymbol{k}}$ has to be paraunitarily diagonalized [42] to find the three physical excitations.
The solution of the coupled magnon-plasmon system can be simplified by first diagonalizing the magnon sector, and then dropping the remaining anomalous magnon-plasmon coupling [35]. The reduced system assumes the form $H^{(2)}\approx\sum_{\boldsymbol{k}}\tilde{\boldsymbol{\Phi}}^{\dagger}_{\boldsymbol{k}}\tilde{H}_{\boldsymbol{k}}\tilde{\boldsymbol{\Phi}}_{\boldsymbol{k}}$, where $\tilde{\boldsymbol{\Phi}}^{\dagger}_{\boldsymbol{k}}=(\alpha^{\dagger}_{\boldsymbol{k}},\beta^{\dagger}_{\boldsymbol{k}},a^{\dagger}_{\boldsymbol{k}})$ contains the magnon normal mode operators $\alpha^{\dagger}_{\boldsymbol{k}}$ and $\beta^{\dagger}_{\boldsymbol{k}}$ that create a magnon with frequency $\omega_{\boldsymbol{k},\pm}=\omega_{\boldsymbol{k}}$. The reduced kernel
$\tilde{H}_{\boldsymbol{k}}=\begin{pmatrix}\hbar\omega_{\boldsymbol{k}}&0&\tilde{G}_{\boldsymbol{k}}\\ 0&\hbar\omega_{\boldsymbol{k}}&\tilde{G}^{\ast}_{\boldsymbol{k}}\\ \tilde{G}^{\ast}_{\boldsymbol{k}}&\tilde{G}_{\boldsymbol{k}}&\hbar\Omega_{\boldsymbol{k}}\end{pmatrix}$ (9)
has an arrow-head matrix structure, and $\tilde{G}_{\boldsymbol{k}}=f_{k}G_{k}\mathrm{e}^{\mathrm{i}\varphi_{\boldsymbol{k}}}$ is the effective magnon-plasmon coupling that—compared to the ferromagnetic case—comes with an enhancement factor $f_{k}>1$, as already pointed out in Ref. [25], see also SM [35].
Figure 3: Magnon-plasmon coupling in an antiferromagnet-metal bilayer. (a) Quasiparticle dispersion, with color indicating the magnon and plasmon character. (b) Thermal Hall conductivity $\kappa_{xy}$ (blue) and spin Nernst conductivity $\alpha_{xy}$ as a function of temperature $T$. Here $\hbar\omega_{0}=1.5\,\mathrm{meV}$ is the exchange-enhanced antiferromagnetic magnon gap.
Given the absence of an external magnetic field, the magnons are spin degenerate and the coupling to the plasmons leads to a spin mixing. We find two anticrossing modes and one (purely magnon-like) mode in the gap [Fig. 3(a)]. Here we have chosen the same parameter values for $J$, $K$, $S$ and $N_{z}$ as for the ferromagnet. The Berry curvature is identically zero, because of an effective time-reversal and inversion symmetry. As a result, $\kappa_{xy}$ is absent, see Fig. 3(b).
Besides heat, magnons carry spin, and a temperature gradient can lead to a transverse spin current, which is quantified by a magnon spin Nernst conductivity [43, 44]
$\alpha_{xy}=\frac{k_{B}}{A\hbar}\sum_{n}\sum_{\boldsymbol{k}}c_{1}(\rho_{n,\boldsymbol{k}})\Omega^{S_{z}}_{n,\boldsymbol{k}}.$ (10)
Here, $c_{1}(x)\equiv x\ln x-(1+x)\ln(1+x)$ and
$\Omega^{S_{z}}_{n,\boldsymbol{k}}=i\sum_{m\neq n}\frac{\langle n|j_{x}^{S}|m\rangle\langle m|\partial_{k_{y}}\tilde{H}_{\boldsymbol{k}}|n\rangle-\langle n|j_{y}^{S}|m\rangle\langle m|\partial_{k_{x}}\tilde{H}_{\boldsymbol{k}}|n\rangle}{(\varepsilon_{\boldsymbol{k},m}-\varepsilon_{\boldsymbol{k},n})^{2}}$ (11)
is the magnon spin Berry curvature, where we defined the spin current $\boldsymbol{j}^{S}=(S_{z}\boldsymbol{v}+\boldsymbol{v}S_{z})/2$ with velocity $\boldsymbol{v}=\hbar^{-1}\partial\tilde{H}_{\boldsymbol{k}}/\partial\boldsymbol{k}$ and spin $S_{z}=\hbar\,\text{Diag}(1,-1,0)$. As shown in Fig. 3(b), we find $\alpha_{xy}$ is non-zero, which is consistent with time-reversal symmetry. We conclude that the coupling to plasmons can be used to engineer spin currents in simple antiferromagnets. Including the anomalous couplings does not qualitatively affect the magnon spin Nernst conductivity at low temperatures, as shown in the SM [35].
Skyrmion crystals.
In both the ferromagnet and the antiferromagnet, we have found a bulk anticrossing, which in the ferromagnet gives rise to nonzero Chern numbers. However, we do not expect to see well-defined chiral edge states because the bare dispersions, $\omega_{k}$ and $\Omega_{k}$, increase monotonously, resulting in a surface spectrum that does not have a band gap to support edge states. Thus, we next explore magnetic systems that naturally feature magnetically active magnon bands with negative curvature in the long-wavelength limit: magnetic skyrmion crystals (SkX). In particular, we focus on counterclockwise rotation (CCW) mode [45] with a negative effective mass.
Topological spin textures have been found in a variety of van der Waals magnets, including metals (Fe3GeTe2 [46], WTe2/Fe3GeTe2 heterostructures [47], and Cr2Ge2Te6 [48]) as well as insulators (CrBr3 [49]). Furthermore, a large Dzyaloshinskii-Moriya interaction was predicted even in a monolayer of Janus van der Waals magnets [50, 51, 52]. Inspired by these experimental observations and computational predictions, we consider skyrmion-hosting van der Waals magnets stacked on graphene and a dielectric spacer. A minimal spin-lattice model is defined on a square lattice as
$H^{(3)}_{\text{magnet}}$ $=$ $-\frac{J}{2}\sum_{\braket{i,j}}\boldsymbol{S}_{i}\cdot\boldsymbol{S}_{j}+\frac{1}{2}\sum_{\braket{i,j}}\boldsymbol{D}_{i,j}\cdot\boldsymbol{S}_{i}\times\boldsymbol{S}_{j}$ (12)
$+$ $g\mu_{B}B_{z}\sum_{i}\boldsymbol{S}_{i}\cdot\hat{\boldsymbol{z}}-K\sum_{i}(S_{i}^{z})^{2}\,$,
with the interfacial Dzyaloshinskii-Moriya interaction $\boldsymbol{D}_{i,j}=D\hat{z}\times(\boldsymbol{r}_{i}-\boldsymbol{r}_{j})/|\boldsymbol{r}_{i}-\boldsymbol{r}_{j}|$.
The SkX becomes the ground state due to the competition between ferromagnetic exchange and Dzyaloshinskii-Moriya interactions at finite magnetic fields, but can be stabilized as a metastable state without external fields [46, 53]. We obtain the classical spin configuration of SkXs by Monte Carlo annealing, which is relaxed via Landau-Lifshitz-Gilbert simulations. Using linear spin wave theory [54], we compute the magnon band structure of the zero-field SkX as shown in Fig. 4(a). We adjust the scaling of the energy and wave vector by assuming $J=10$ meV, $D=1$ meV, $K_{z}=0.025$ meV, and the size of magnetic unit cell L=100L=100~nm [35].
Since the CCW mode gives rise to the rotation of the in-plane magnetization [55, 35], the magnon-plasmon coupling induces a winding in hybridized wave functions with a topological gap. In addition, the topological spin textures of SkXs support the nontrivial magnon band topology [56, 57, 58]. While the CCW mode is characterized by C=1C=-1, the magnon-plasmon coupling induces $C=+1$ below the hybridization gap. Thus, chiral edge modes of magnons and magnon-plasmon hybrids propagate in opposite directions, which might result in backscattering. However, a significant mismatch in energy and wave vector of these edge modes suppresses scattering among them. Consequently, both edge modes should remain robust, although they are not topologically protected.
We construct the tight-binding model of magnon-plasmon hybrids in SkXs by fitting the dispersion of plasmons with quadratic terms [35]. Employing the renormalization technique for the semi-infinite geometry [59, 60], we compute the local density of states (LDOS) of magnon-plasmon hybrids at an edge of the semi-infinite skyrmion lattice as shown in Fig. 4(b) and (c), respectively. The damping rate is set at $\kappa_{\textrm{mag}}=\kappa_{\textrm{pl}}=4\textrm{ MHz}\approx 10^{-4}\omega_{\textrm{res}}$ with the resonance frequency $\omega_{\textrm{res}}=39$ GHz. Since the coupling strength between the CCW and plasmon modes is estimated as $G\sim 10^{-2}\omega_{\textrm{res}}$ [35], the strong coupling regime is realized in this setup. Figure 4(b) shows the LDOS over the one-dimensional first Brillouin zone of semi-infinite SkXs. The magnonic chiral edge mode is found inside the topological gap of the CCW mode, while it does not exhibit a clear signature of the anticrossing with plasmons. With a high resolution in energy and wave vector in Fig. 4(c), we obtain the magnon-plasmon edge state that connects the lower magnon branch with the upper plasmon branch. The LDOS of plasmon bands is not visible in Fig. 4(c) due to the significant difference in their effective masses. The group velocity of the magnon-plasmon edge mode is opposite to the magnonic edge mode above the CCW mode, which is consistent with the Chern number calculation.
Figure 4: Topological magnon-plasmon polariton in a heterostructure of SkXs and graphene. (a) Three-dimensional magnon band structures of the metastable zero-field SkX. The energy and wave vector are rescaled from the dimensionless spin-lattice model as detailed in the SM [35]. (b,c) Local density of states (LDOS) of magnon-plasmon hybrids at an edge of the semi-infinite skyrmion lattice.
We also investigate the effect of damping in the magnon-plasmon edge state [35]. As the damping rate of magnon modes increases, it becomes harder to identify the magnon-plasmon edge state due to the broadening of the bulk spectrum. We find that the chiral edge state is robust up to $\kappa_{\textrm{mag}}=10^{-3}\omega_{\textrm{mag}}$. In contrast, the magnon-plasmon edge state is less sensitive to the damping rate of plasmons, indicating the dominant magnonic contributions in the LDOS. The damping rate of plasmons was recently measured to be $3.1\times 10^{-4}\omega_{\textrm{pl}}$ in the infrared light [61]. In addition, the damping rate of magnons in magnetic insulators is approximated as $\kappa_{\textrm{mag}}=\alpha\omega_{\textrm{mag}}$, where the lowest value of the Gilbert damping $\alpha$ is known as $\alpha=3\times 10^{-5}$ at room temperature in yttrium iron garnet [62]. Thus, sufficiently small magnon and plasmon damping rates are attainable at cryogenic conditions.
The hybrid nature of the chiral edge state is advantageous for experimental observations. Even though observing the chiral edge state has been challenging in topological magnonics, they have been successfully observed in topological plasmonics and photonics [63, 64]. For example, the plasmonic chiral edge state can be excited by gold antennas [65] and detected by the near-field microscropy [63, 61] and the transmission spectrum [66]. In addition, magnonic components can be measured by the NV center magnetometry [67, 68] and near-field Brillouin light scattering [69].
Discussion and conclusion.
We have identified topologically nontrivial magnon-plasmon hybrids and chiral magnon-plasmon edge states in effectively two-dimensional stacks of van der Waals layers. The resulting nontrivial band geometry was shown to lead to transverse heat and spin transport and to chiral edge magnon-plasmon hybrids. Our effective theory and, in particular, the lattice regularization required to calculate the edge spectrum in Fig. 4(b,c), rely on the quasiparticle approximation and neglect the Landau damping in the electron-hole continuum at larger momenta. For open boundaries, the magnon-plasmon gap is filled with a continuum of states. We expect that the hybridization of single-particle states with the continuum will lead to breakdown effects similar to those discussed in Ref. 70 for interacting topological magnon insulators and in Ref. 71 for a generic bosonic toy model. However, as long as the energy scale of these effects is smaller than the topological bulk gap, our effective treatment captures the relevant single-particle physics, and an effective non-Hermitian description as in Ref. 72 could be pursued. To avoid the continuum altogether, one could use intrinsically undamped plasmon modes in narrow electron bands, e.g., in Moire graphene bilayers [73].
For the coupling of magnons to plasmons, we have relied on the polaritonic component and magnetic dipole coupling, a small effect but sufficient for a proof of principle. Importantly, magnon-plasmon coupling can occur in a single material with spin-orbit coupling [25], which also gives rise to the crucial phase winding necessary for the Berry curvature. The range of possible material platforms for the realization of topological magnon plasmons is thus not limited to the stacked layers considered here. The case of plasmons hybridizing with magnon continua, as studied in Refs. 24, 28, could also be attractive for the realization of chiral spin-plasma edge modes. A critical challenge will be to enhance the magnon-plasmon cooperativity by reducing the quasiparticle damping and increasing the coupling to stabilize wide topological gaps supporting chiral modes, which may find application in nonreciprocal devices such as circulators.
Note added.
After submission of our work, we became aware of related studies on topological plasmon–magnon modes in magnetic heterostructures of a Rashba electron gas and a magnetic system [74].
Acknowledgements.
Acknowledgments. We thank Johannes Knolle, Daniel Malz, Peng Rao, and António Costa for stimulating discussions. This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - Project No. 504261060. T. H. is supported by JSPS KAKENHI Grant Number JP23K13064 and Aoyama Gakuin University Research Institute “Early Eagle” grant program for promotion of research by early career researchers. P. G. acknowledges financial support from the Alexander von Humboldt postdoctoral fellowship.